OpenStudy (anonymous):

Algebra 1 Homework Help? Please?

OpenStudy (anonymous):

I'm really bad at Algebra, and I need some help figuring out some problems. For example: y = 4/8x - 45

OpenStudy (anonymous):

you're supposed to graph this don't you?

OpenStudy (anonymous):

There's no graphs available for this question. No.

OpenStudy (anonymous):

What he is saying is there is nothing to solve here

OpenStudy (anonymous):

because it's in y=mx+b form

OpenStudy (anonymous):

What are you trying to find?

OpenStudy (anonymous):

Oh, my apologies. That's suppose to be a 9, not an 8. y = 4/9x - 45 But there's multiple choice for it, which is: a. –5 b. 0.2 c. 5 d. –0.2

OpenStudy (anonymous):

is there a question that goes along with this problem?

OpenStudy (anonymous):

There's two of them, I was hoping to have an example of how to work it so I could attempt to figure out the second. But here:

OpenStudy (anonymous):

Ah I see now. You see this is just an equation for a graph it really doesn't ask you to solve anything but when you show the question now I now what to solve for.

OpenStudy (anonymous):

Ohh sorry i was confused since you didn't write parentheses like thisy = 4/(9x - 45)

OpenStudy (anonymous):

So you know that the denomintaor can't be zero right? Never. because it doesn't exist so what would make the denominator zero?

OpenStudy (anonymous):

By having a common denominator? I'm not sure, my teacher did a horrid job in trying to explain it to me, which is why I'm here, so I apologize if I am far off.

OpenStudy (anonymous):

NO if you substituted a number for x. Which one would give you zero in the denominator?

OpenStudy (anonymous):

you have to find something thhat makes the denominator zero, so ore example y = 4/(9x - 45) it would be 5, since 9*5 is 45, so then 45-45=0

OpenStudy (anonymous):

for*

OpenStudy (anonymous):

Oh, okay! That makes perfect sense actually!

OpenStudy (anonymous):

So you understand? Can you do the second one?

OpenStudy (anonymous):

So, the second one would be 24?

OpenStudy (anonymous):

-24

OpenStudy (anonymous):

right -24

OpenStudy (anonymous):

Where did the negative come from?

OpenStudy (anonymous):

what's -24 *2

OpenStudy (anonymous):

-48

OpenStudy (anonymous):

and what's -48 +48

OpenStudy (anonymous):

0

OpenStudy (anonymous):

and that's why it's -24.

OpenStudy (anonymous):

Oh, okay now I understand. And I have another question I need assistance, this one does involve graphing.

OpenStudy (anonymous):

ok not a problem

OpenStudy (anonymous):

I'm not sure on how to work it (the rest of the answer selection is in picture 2 (named 3 by mistake)

OpenStudy (anonymous):

ok this is pretty much the same deal. What number for x will give you a zero in the denominator?

OpenStudy (anonymous):

2? I'm lost.

OpenStudy (anonymous):

no the 2 isn't in the denominator at all. that -2 is just the y-intercept. just pay attention to this \[\frac{ 1 }{ x-4 }\] what number if you substitute it for x will give you zero on the bottom

OpenStudy (anonymous):

Either a positive 4, or a 3?..

OpenStudy (anonymous):

ok, well if you aren't sure go ahead and plug them in what is 4-4 and what is 3-4

OpenStudy (anonymous):

4-4 = 0 3-4 = -1

OpenStudy (anonymous):

And the reason I even brought up the three, is because my teacher said to always calculate the number on top with the one on the bottom, which, in fact, has not been working well.

OpenStudy (anonymous):

that might come in handy later but we'll cross that bridge when we come to it

OpenStudy (anonymous):

anyway now the question you need to ask yourself is whether the asymptote crosses at y=4 or x=4. Any ideas?

OpenStudy (anonymous):

Possibly x=4?

OpenStudy (anonymous):

yes that's correct. It's a vertical asymptote not a horizontal if it crossed at y=4 it would look something like this